Solve each problem and show the solutions.
1.find the sum of the first 20 integers
2.find the sum of the first 30 even integers
3.find the sum of the integers from 4 to 128

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

1. This is n(n+1)/2=20*21/2=210
2. This is n(n+1)=30*31=930
3. sum from 4 to 128
Sum of first 128 is 128*129/2=8256
sum of first 3 is 6
the difference is the answer, or 8250.

Answer by ikleyn(52788)

.

For your info:

            In  Math,  there is no such a notion as  " first  20  integers "     < < < - - - = = =    because it  MAKES  no  SENSE.

            There is a notion  " first  20  positive integers ",  INSTEAD,  if it is what you mean,  in reality.